# GeoAware-VLA: Implicit Geometry Aware Vision-Language-Action Model

### v1

@@ Abstract
Vision-Language-Action (VLA) models often fail to generalize to novel camera viewpoints, a limitation stemming from their difficulty in inferring robust 3D geometry from 2D images. We introduce GeoAware-VLA, a simple yet effective approach that enhances viewpoint invariance by integrating strong geometric priors into the vision backbone. Instead of training a visual encoder or relying on explicit 3D data, we leverage a frozen, pretrained geometric vision model as a feature extractor. A trainable projection layer then adapts these geometrically-rich features for the policy decoder, relieving it of the burden of learning 3D consistency from scratch. Through extensive evaluations on LIBERO benchmark subsets, we show GeoAware-VLA achieves substantial improvements in zero-shot generalization to novel camera poses, boosting success rates by over 2x in simulation. Crucially, these benefits translate to the physical world; our model shows a significant performance gain on a real robot, especially when evaluated from unseen camera angles. Our approach proves effective across both continuous and discrete action spaces, highlighting that robust geometric grounding is a key component for creating more generalizable robotic agents.
@@ I Introduction
The development of general-purpose agents capable of performing diverse manipulation tasks in unstructured environments is a central goal in robotics. Vision-Language-Action (VLA) model is a popular paradigm, which learns to map visual observations and natural language instructions directly to robot actions. Most methods achieve strong performance within their training domains [1, 2, 3], but often fail to generalize beyond them, struggling even with minor changes in camera viewpoints [4, 5]. This limitation arises from the difficulty of inferring a consistent 3D world model from 2D visual inputs, a prerequisite for reliable manipulation and spatial awareness.
Prior work has explored two main strategies to address this challenge. One approach is to incorporate explicit 3D representations, such as point clouds, into the policy’s observation space [6, 7, 8]. While effective, these methods often require depth sensors and introduce significant computational overhead for constructing and processing explicit 3D structures. An alternative lies in implicitly encouraging the model to learn geometrically consistent features without reconstructing the scene in 3D. This is often achieved by leveraging multi-view data or data augmentation during training to produce a view-invariant policy [9, 10].
@@
However, these implicit methods also have their own drawbacks. Disentanglement techniques often rely on curated multi-view datasets for effective training, while augmentation-based methods are fundamentally limited by the computational cost of producing novel views and the distribution of views from which they are sampled. These challenges can be addressed by leveraging a powerful, pretrained geometric foundation models that have already distilled 3D understanding from vast datasets. For instance, the Visual Geometry Grounded Transformer (VGGT) [11] is explicitly optimized to distill a rich 3D representation of the world, having been trained to infer key geometric aspects such as camera parameters, multi-view depth, dense point clouds, and point tracking.
Built on these insights, we propose GeoAware-VLA, a simple yet highly effective modification to a standard VLA architecture that significantly enhances robustness to novel camera viewpoints. Our core hypothesis is that a policy’s ability to generalize across views is fundamentally tied to the geometric acuity of its visual encoder. We therefore replace the standard image encoder with the powerful VGGT backbone, using it as a frozen feature extractor. By building upon features that are already inherently consistent across different perspectives, our policy is relieved of the burden of learning 3D geometry from scratch. We need only append a lightweight, trainable projection layer to map these potent features into the latent space of the BAKU policy decoder [1]. As our experiments demonstrate, the integration of a powerful geometric prior results in substantial improvements in zero shot generalization to novel camera viewpoints as shown in Figure 1.
@@ II-A Imitation Learning
Modern robotic learning is rooted in Imitation Learning (IL), where early methods like Behavioral Cloning (BC) [12] struggled with compounding errors from distribution shift. This motivated the development of more robust visuomotor policies, with generative methods like Diffusion Policy [13] and energy-based models like Implicit Behavioral Cloning (IBC) [14] learning complex action distributions to enhance control. Building on this robustness, the field’s focus shifted towards creating large-scale, generalist agents. Influential models like DeepMind’s Gato [15] and Octo [16] demonstrated that a single, large transformer pre-trained on diverse, multi-robot datasets could establish a powerful and adaptable motor control foundation. This trajectory culminated in the current paradigm of VLA models, which unify perception, language, and control by leveraging knowledge from web-scale models. This includes a series of influential works such as RT-1 [17] and its successors [18, 19], alongside powerful open-source models like OpenVLA [2, 20] and $\pi$ -[21, 22], which have collectively pushed the boundaries of generalization and performance. Our work builds on the BAKU architecture [1], a lightweight VLA that stands out for its efficient design choices, including multi-sensory observations and action chunking.
@@ II-B Vision Encoders in Robotic Learning
The vision backbone, which processes raw pixel inputs, is a critical component of VLA models. Foundational approaches often used semantic encoders like ResNet [23], and later enhanced with FiLM layers to incorporate conditioning from language or other inputs [24].
A more recent paradigm involves models like SigLIP [25] and its successor SigLIP 2 [26], which are pre-trained on vast web-scale image-text pairs. This training method creates representations that are naturally aligned with language, making them highly effective for VLA tasks. However, the primary focus of these backbones remains semantic - identifying what objects are, rather than their precise 3D geometry.
Alongside the development of semantic encoders for policy learning, a distinct class of powerful geometric vision models has emerged from the computer vision community. Architectures such as DUSt3R [27], MUST3R [28], and VGGT [11] are systems designed specifically for geometric inference from multi-view imagery, inferring properties like camera poses and dense depth maps. While highly effective for 3D reconstruction, these models were not intended as general-purpose backbones for policy learning and, consequently, their use in robotics has been limited. To our knowledge, such geometric models have not been implicitly integrated into VLA models before. Our work provides an initial exploration in this direction, investigating how to bridge this gap by leveraging the implicit 3D understanding from these geometric models to improve robotic policies.
@@ II-C View-point robust robotic policies
To achieve more viewpoint-robust policies, various methods have been explored. One approach involves augmenting training data with images rendered from simulators [9, 29]. While effective, this introduces the challenge of sim-to-real transfer. Another promising way is utilizing novel view synthesis [10, 30]. Unlike simulator-based methods, these models can operate directly on real-world data, circumventing the sim-to-real gap. However, these models may still perform poorly when the new viewpoint is drastically different from the original view [10].
Other works, like Act3D [31] and 3D Diffuser Actor [32], utilize explicit 3D representations such as point clouds or voxel grids [33, 34, 8, 6]. These methods are robust to viewpoint changes but often require reliable depth data and camera calibration, which can be scarce in large-scale datasets like [19, 35, 36]. In contrast to these methods, our approach integrates strong geometric priors, which relieves the policy of the burden of learning 3D consistency from scratch and avoids the need for explicit depth or calibration information.
@@ IV Methodology
We propose GeoAware-VLA, a novel approach to enhance a standard VLA framework. The core of our method is the replacement of the trainable vision encoder with a frozen, Geometrically-Aware one. To make the rich, multi-layer features from the VGGT backbone usable by the policy network, we introduce a trainable vision projection layer. The remainder of the policy architecture adapts and builds upon elements from [1]. The overall architecture of our model is illustrated in Figure 2
@@ IV-A2 Language and Proprioceptive Encoders
  • Language Encoder: A task instruction string $l_{t}$ is first encoded using a pretrained sentence transformer [37], yielding an embedding of dimension $D_{\text{lang\_emb}}$. An MLP then projects this embedding into the common representation space, producing $\mathbf{z}_{\text{lang}}\in\mathbb{R}^{D_{\text{repr}}}$.
@@ IV-C Action Heads
Table I: Success Rates (%) on Libero Subsets. Models marked with * are results taken from [20].
@@
Diffusion Policy* [13] 78.3 - 92.5 - 68.3 - 50.5 - 72.4 -
Octo* [16] 78.9 - 85.7 - 84.6 - 51.1 - 75.1 -
OpenVLA-OFT [20] 98.0 90.0 98.0 15.7 98.0 81.7 91.0 13.7 96.3 50.2
BAKU [1] 94.0 18.0 100.0 49.3 96.0 80.7 87.0 3.7 94.2 37.9
VQ-BeT [38] 94.0 41.3 100.0 38.3 98.0 83.0 93.0 3.0 96.3 41.4
@@
OpenVLA-OFT [20] 59.3 47 44.5 50.2
BAKU [1] 41.5 31.5 40.8 37.9
VQ-BeT [38] 44.0 43.8 36.5 41.4
@@ V-A Simulation Benchmarks
The LIBERO benchmark [35] assesses knowledge transfer in multitask and lifelong robot learning for manipulation tasks. It is composed of four task suites: LIBERO-Spatial, which tests generalization to novel spatial layouts while keeping task and object types fixed; LIBERO-Goal, which introduces new tasks while maintaining the same objects and layouts; LIBERO-Object, which evaluates performance with Novel object types under the same tasks and layouts; and LIBERO-Long, which presents a broader variety of objects, layouts, and backgrounds. Each suite consists of 10 tasks, with 50 human demonstrations provided per task for fine-tuning. We evaluate models trained on Libero’s suites on original and novel viewpoints inspired by [6] both of which are shown in Figure 3.
@@ V-C Baselines
We compare our GeoAware-VLA model against three strong baselines: two derived from [1] architecture and OpenVLA-OFT [20], which we include to test performance with a larger, pretrained model.
@@ V-D Training
All models were implemented in PyTorch and trained on a single NVIDIA A100 GPU with 40GB of VRAM. For experiments on the LIBERO benchmark, we trained a separate policy for each of the four task suites. Each model was trained for 150,000 steps using the AdamW optimizer. For our real-world evaluation, policies were trained for 50,000 steps on our custom dataset. Across all experiments, we used a batch size of 64 and provided the policy with observations from two camera views.
@@ VI Results and Discussions
We evaluate GeoAware-VLA to answer three key questions:
@@ VI-A A Geometric Backbone is a Strong In-Distribution Performer
This result supports our hypothesis that a lightweight, trainable projection layer is sufficient to adapt the geometrically-rich features from a frozen backbone for the policy. It demonstrates that relying on a geometric prior provides an effective foundation for policy learning without compromising in-distribution performance.
@@ VI-D Ablation on VGGT Layer Selection
As detailed in Table III, the default configuration using evenly spaced layers achieves the best overall performance. While using all layers provides a small improvement in novel view generalization, it results in a slight drop in in-distribution performance and is more computationally expensive. It is worth noting that even the weakest configuration using the last 4 layer significantly outperforms the next-best baseline on novel views. Despite its overall performance drop, its 34.3% success rate on unseen views is still more than double that of [20].

### v3

@@ Abstract
Vision-Language-Action (VLA) models often fail to generalize to novel camera viewpoints, a limitation stemming from their difficulty in inferring robust 3D geometry from 2D images. We introduce GeoAware-VLA, a simple yet effective approach that enhances viewpoint invariance by integrating strong geometric priors into the vision backbone. Instead of training a visual encoder or relying on explicit 3D data, we leverage a frozen, pretrained geometric vision model as a feature extractor. A trainable projection layer then adapts these geometrically-rich features for the policy decoder, relieving it of the burden of learning 3D consistency from scratch. Through extensive evaluations on LIBERO benchmark subsets, we show GeoAware-VLA achieves substantial improvements in zero-shot generalization to novel camera poses, boosting success rates by over 2x in simulation. Crucially, these benefits translate to the physical world; our model shows a significant performance gain on a real robot, especially when evaluated from unseen camera angles. Our approach proves effective across both continuous and discrete action spaces, highlighting that robust geometric grounding is a key component for creating more generalizable robotic agents. Project Page https://alisharey.github.io/GeoAware-VLA/
@@ I Introduction
The development of general-purpose agents capable of performing diverse manipulation tasks in unstructured environments is a central goal in robotics. Vision-Language-Action (VLA) model is a popular paradigm, which learns to map visual observations and natural language instructions directly to robot actions. Most methods achieve strong performance within their training domains [haldar2024baku, kim2024openvla, zitkovich2023rt], but often fail to generalize beyond them, struggling even with minor changes in camera viewpoints [simpler, xie2024decomposing]. This limitation arises from the difficulty of inferring a consistent 3D world model from 2D visual inputs, a prerequisite for reliable manipulation and spatial awareness.
Prior work has explored two main strategies to address this challenge. One approach is to incorporate explicit 3D representations, such as point clouds, into the policy’s observation space [wilcox2025adapt3r, yang2025fp3, ze20243d]. While effective, these methods often require depth sensors and introduce significant computational overhead for constructing and processing explicit 3D structures. An alternative lies in implicitly encouraging the model to learn geometrically consistent features without reconstructing the scene in 3D. This is often achieved by leveraging multi-view data or data augmentation during training to produce a view-invariant policy [pang2025learning, tian2024view].
@@
However, these implicit methods also have their own drawbacks. Disentanglement techniques often rely on curated multi-view datasets for effective training, while augmentation-based methods are fundamentally limited by the computational cost of producing novel views and the distribution of views from which they are sampled. These challenges can be addressed by leveraging a powerful, pretrained geometric foundation models that have already distilled 3D understanding from vast datasets. For instance, the Visual Geometry Grounded Transformer (VGGT) [wang2025vggt] is explicitly optimized to distill a rich 3D representation of the world, having been trained to infer key geometric aspects such as camera parameters, multi-view depth, dense point clouds, and point tracking.
Built on these insights, we propose GeoAware-VLA, a simple yet highly effective modification to a standard VLA architecture that significantly enhances robustness to novel camera viewpoints. Our core hypothesis is that a policy’s ability to generalize across views is fundamentally tied to the geometric acuity of its visual encoder. We therefore replace the standard image encoder with the powerful VGGT backbone, using it as a frozen feature extractor. By building upon features that are already inherently consistent across different perspectives, our policy is relieved of the burden of learning 3D geometry from scratch. We need only append a lightweight, trainable projection layer to map these potent features into the latent space of the BAKU policy decoder [haldar2024baku]. As our experiments demonstrate, the integration of a powerful geometric prior results in substantial improvements in zero shot generalization to novel camera viewpoints as shown in Figure 1.
@@ II-A Imitation Learning
Modern robotic learning is rooted in Imitation Learning (IL), where early methods like Behavioral Cloning (BC) [pomerleau1988alvinn] struggled with compounding errors from distribution shift. This motivated the development of more robust visuomotor policies, with generative methods like Diffusion Policy [chi2023diffusion] and energy-based models like Implicit Behavioral Cloning (IBC) [florence2022implicit] learning complex action distributions to enhance control. Building on this robustness, the field’s focus shifted towards creating large-scale, generalist agents. Influential models like DeepMind’s Gato [reedgeneralist] and Octo [ghosh2024octo] demonstrated that a single, large transformer pre-trained on diverse, multi-robot datasets could establish a powerful and adaptable motor control foundation. This trajectory culminated in the current paradigm of VLA models, which unify perception, language, and control by leveraging knowledge from web-scale models. This includes a series of influential works such as RT-1 [RT-1] and its successors [RT-2, RT-X], alongside powerful open-source models like OpenVLA [kim2024openvla, kim2025fine] and $\pi$ -[black2410pi0, intelligence2504pi05], which have collectively pushed the boundaries of generalization and performance. Our work builds on the BAKU architecture [haldar2024baku], a lightweight VLA that stands out for its efficient design choices, including multi-sensory observations and action chunking.
@@ II-B Vision Encoders in Robotic Learning
The vision backbone, which processes raw pixel inputs, is a critical component of VLA models. Foundational approaches often used semantic encoders like ResNet [he2016deep], and later enhanced with FiLM layers to incorporate conditioning from language or other inputs [perez2018film].
A more recent paradigm involves models like SigLIP [zhai2023sigmoid] and its successor SigLIP 2 [tschannen2025siglip], which are pre-trained on vast web-scale image-text pairs. This training method creates representations that are naturally aligned with language, making them highly effective for VLA tasks. However, the primary focus of these backbones remains semantic - identifying what objects are, rather than their precise 3D geometry.
Alongside the development of semantic encoders for policy learning, a distinct class of powerful geometric vision models has emerged from the computer vision community. Architectures such as DUSt3R [wang2024dust3r], MUST3R [cabon2025must3r], and VGGT [wang2025vggt] are systems designed specifically for geometric inference from multi-view imagery, inferring properties like camera poses and dense depth maps. While highly effective for 3D reconstruction, these models were not intended as general-purpose backbones for policy learning and, consequently, their use in robotics has been limited. To our knowledge, such geometric models have not been implicitly integrated into VLA models before. Our work provides an initial exploration in this direction, investigating how to bridge this gap by leveraging the implicit 3D understanding from these geometric models to improve robotic policies.
@@ II-C View-point robust robotic policies
To achieve more viewpoint-robust policies, various methods have been explored. One approach involves augmenting training data with images rendered from simulators [pang2025learning, seo2023multi]. While effective, this introduces the challenge of sim-to-real transfer. Another promising way is utilizing novel view synthesis [tian2024view, yang2025mobi]. Unlike simulator-based methods, these models can operate directly on real-world data, circumventing the sim-to-real gap. However, these models may still perform poorly when the new viewpoint is drastically different from the original view [tian2024view].
Other works, like Act3D [gervet2023act3d] and 3D Diffuser Actor [ke20243ddiffuser], utilize explicit 3D representations such as point clouds or voxel grids [shridhar2023perceiver, zhulearning, ze20243d, wilcox2025adapt3r]. These methods are robust to viewpoint changes but often require reliable depth data and camera calibration, which can be scarce in large-scale datasets like [RT-X, libero, walke2023bridgedata]. In contrast to these methods, our approach integrates strong geometric priors, which relieves the policy of the burden of learning 3D consistency from scratch and avoids the need for explicit depth or calibration information.
@@ IV Methodology
We propose GeoAware-VLA, a novel approach to enhance a standard VLA framework. The core of our method is the replacement of the trainable vision encoder with a frozen, Geometrically-Aware one. To make the rich, multi-layer features from the VGGT backbone usable by the policy network, we introduce a trainable vision projection layer. The remainder of the policy architecture adapts and builds upon elements from [haldar2024baku]. The overall architecture of our model is illustrated in Figure 2
@@ IV-A2 Language and Proprioceptive Encoders
  • Language Encoder: A task instruction string $l_{t}$ is first encoded using a pretrained sentence transformer [reimers-2019-sentence-bert], yielding an embedding of dimension $D_{\text{lang\_emb}}$. An MLP then projects this embedding into the common representation space, producing $\mathbf{z}_{\text{lang}}\in\mathbb{R}^{D_{\text{repr}}}$.
@@ IV-C Action Heads
Table I: Success Rates (%) on Libero Subsets. Models marked with * are results taken from [kim2025fine].
@@
Diffusion Policy* [chi2023diffusion] 78.3 - 92.5 - 68.3 - 50.5 - 72.4 -
Octo* [ghosh2024octo] 78.9 - 85.7 - 84.6 - 51.1 - 75.1 -
OpenVLA-OFT [kim2025fine] 98.0 90.0 98.0 15.7 98.0 81.7 91.0 13.7 96.3 50.2
BAKU [haldar2024baku] 94.0 18.0 100.0 49.3 96.0 80.7 87.0 3.7 94.2 37.9
VQ-BeT [lee2024behavior] 94.0 41.3 100.0 38.3 98.0 83.0 93.0 3.0 96.3 41.4
@@
OpenVLA-OFT [kim2025fine] 59.3 47 44.5 50.2
BAKU [haldar2024baku] 41.5 31.5 40.8 37.9
VQ-BeT [lee2024behavior] 44.0 43.8 36.5 41.4
@@ V-A Simulation Benchmarks
The LIBERO benchmark [libero] assesses knowledge transfer in multitask and lifelong robot learning for manipulation tasks. It is composed of four task suites: LIBERO-Spatial, which tests generalization to novel spatial layouts while keeping task and object types fixed; LIBERO-Goal, which introduces new tasks while maintaining the same objects and layouts; LIBERO-Object, which evaluates performance with Novel object types under the same tasks and layouts; and LIBERO-Long, which presents a broader variety of objects, layouts, and backgrounds. Each suite consists of 10 tasks, with 50 human demonstrations provided per task for fine-tuning. We evaluate models trained on Libero’s suites on original and novel viewpoints inspired by [wilcox2025adapt3r] both of which are shown in Figure 3.
@@ V-C Baselines
We compare our GeoAware-VLA model against three strong baselines: two derived from [haldar2024baku] architecture and OpenVLA-OFT [kim2025fine], which we include to test performance with a larger, pretrained model.
@@ V-D Training
All models were implemented in PyTorch and trained on a single NVIDIA A100 GPU with 40GB of VRAM. For experiments on the LIBERO benchmark, we trained a separate policy for each of the four task suites. Each model was trained for 150,000 steps using the AdamW optimizer. For our real-world evaluation, policies were trained for 50,000 steps on our custom dataset. Across all experiments, we used a batch size of 64 and provided the policy with observations from two camera views. GeoAware variants were trained with input images resized to $126\times 126$ pixels, whereas the BAKU variants used $128\times 128$ pixels.
@@ VI Results and Discussions
We evaluate GeoAware-VLA to answer four key questions:
@@ VI-A A Geometric Backbone is a Strong In-Distribution Performer
This result supports our hypothesis that a lightweight, trainable projection layer is sufficient to adapt the geometrically-rich features from a frozen backbone for the policy.
@@ VI-D Ablation on VGGT Layer Selection
As detailed in Table III, the default configuration using evenly spaced layers achieves the best overall performance. While using all layers provides a small improvement in novel view generalization, it results in a slight drop in in-distribution performance and is more computationally expensive. It is worth noting that even the weakest configuration using the last 4 layer significantly outperforms the next-best baseline on novel views. Despite its overall performance drop, its 34.3% success rate on unseen views is still more than double that of [kim2025fine].